OpenStudy (anonymous):

Find f(a + h) - f(a) for the function. Simplify your answer. f(x)=x^2-9x+3. I'm not exactly sure how to go about this...

OpenStudy (anonymous):

first understand what \[f(a+h)\] means. it means replace x everywhere you see it by \[a+h\] so you get \[f(a+h)=(a+h)^2-9(a+h)+3\]

OpenStudy (anonymous):

and of course \[f(a)=a^2-9a+3\]

OpenStudy (anonymous):

so \[f(a+h)-f(a)=(a+h)^2-9(a+h)+3-(a^2-9a+3)\] and it is algebra from here on in

OpenStudy (anonymous):

a bunch of algebra in fact \[a^2+2ah+h^2-9a-9h+3-a^2+9a-3\] now the usual miracle occurs, everything without an h in it goes away and you get \[2ah-9h+h^2\]

OpenStudy (anonymous):

I see now. I was forgetting how to start this type of problem, but right after your first post I realized what I had to do. Thanks a ton!